Wersja testu A 16 września 2002r.

1. a. T b. N c. T d. T 2. a. N b. T c. T d. N 3. a. N b. N c. T d. N 4. a. T b. N c. T d. N 5. a. T b. N c. N d. T 6. a. N b. N c. N d. N 7. a. T b. T c. T d. N 8. a. T b. T c. N d. N 9. a. N b. N c. N d. T 10. a. T b. N c. T d. N 11. a. T b. N c. T d. N 12. a. T b. T c. T d. N 13. a. T b. T c. N d. N 14. a. T b. N c. N d. N 15. a. N b. T c. T d. T 16. a. N b. N c. N d. T 17. a. T b. T c. N d. N 18. a. T b. T c. N d. T 19. a. N b. T c. N d. T 20. a. T b. N c. N d. T